Solve each system by addition.
2x=2-y)+(22x+36y=356)
i got a decimal for y and i didnt think it was right.


Answer by richwmiller(17219)

so what did you do and what did you get?
2x=2-y
changing it the the same order as the second equation
2x+y=2
multiply by -11
-22x+-11y=-22
22x+36y=356
add
25y=334
y=334/25 x=-142/25
They are either fractions over 25 or decimals